OpenStudy (anonymous):

is log(8x)-log(x^2-1)=log(3) be the same as 8x-(x^2-1)=3

jimthompson5910 (jim_thompson5910):

no it is not

jimthompson5910 (jim_thompson5910):

you need to combine the logs on the left side using this formula log(A) - log(B) = log(A/B)

OpenStudy (anonymous):

(sorry I don't think it will make a difference but all the logs are base 2)

jimthompson5910 (jim_thompson5910):

no that doesn't affect the answer

OpenStudy (anonymous):

ok... so what would this be in this case? would that rule apply if it was log()+log()?

jimthompson5910 (jim_thompson5910):

if you add 2 logs of the same base, you use log(A)+log(B) = log(A*B)

OpenStudy (anonymous):

sorry to ask this, but in what case would that rule apply then?

jimthompson5910 (jim_thompson5910):

which rule

OpenStudy (anonymous):

log() log()=log() is the same as () ()=() (I thought the symbols inside were + or -)

OpenStudy (anonymous):

I originally thought it was log()+log()=log() is the same as ()+()=()

jimthompson5910 (jim_thompson5910):

no the log rules are given here http://www.purplemath.com/modules/logrules.htm

jimthompson5910 (jim_thompson5910):

it's handy to have them memorized

jimthompson5910 (jim_thompson5910):

once you have log(A) = log(B), you can say A = B

OpenStudy (anonymous):

Ah I see.

jimthompson5910 (jim_thompson5910):

that's why you combine the logs on the left side to have log = log

OpenStudy (anonymous):

so I don't have to redo log(-3x-1)=log(-4x-6)?

jimthompson5910 (jim_thompson5910):

no because that's in the form log(A) = log(B) A = -3x-1 B = -4x-6

OpenStudy (anonymous):

ah I see. so... sorry to ask this, but can you help me through with these types of problems? I have log(8x/x^2-1)=log(3)

jimthompson5910 (jim_thompson5910):

so you can now say \[\Large \frac{8x}{x^2-1} = 3\]

OpenStudy (anonymous):

would the x and x^2 cancel so it becomes log(8/x-1)?

jimthompson5910 (jim_thompson5910):

no you can't do that

OpenStudy (anonymous):

Ok sorry I didn't notice the log() removed.

jimthompson5910 (jim_thompson5910):

also, you can't cancel the x's like that either

OpenStudy (anonymous):

multiply both sides by x^2-1 to get 8x=3x^2-3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

get everything to one side, then use the quadratic formula

jimthompson5910 (jim_thompson5910):

you could factor, but factoring doesn't always work

OpenStudy (anonymous):

so then zero out one side 3x^2-8x-3=0 then factor it? (3x+1)(x-3)

jimthompson5910 (jim_thompson5910):

that works too

OpenStudy (anonymous):

then, the options for an answer would be 0, -1/3, and 3?

OpenStudy (anonymous):

just need to check which one works and take out the option for 0?

jimthompson5910 (jim_thompson5910):

how are you getting 0?

OpenStudy (anonymous):

sorry IDK must've been a mistake. but, it should be log(8(3))-log(9-1)=log(3) log(24)-log(8)=log(3) log(24/8)=log(3) 24/8=3

jimthompson5910 (jim_thompson5910):

so 3 is definitely a solution, check -1/3 as well

OpenStudy (anonymous):

sorry it took a while

jimthompson5910 (jim_thompson5910):

its fine, take all the time you need

OpenStudy (anonymous):

log(8(-1/3))-log(-1/3)^2-1)=log(3) log(-8/3)-log(1/9-1)=log(3) log(-8/3/1/9-1)=log(3) (-8/3)/(1/9)-1=log(3) (-8/3)/(-8/9)=3 3=3 so both 3 and -1/3 are answers

OpenStudy (anonymous):

did I do it right?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Thank you! my mind was blown when I realized how simple it was! Thank you thank you thank you.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I have one more question... I need someone to check it to make sure it's correct

OpenStudy (anonymous):

3^x=17 solve for the variable

OpenStudy (anonymous):

this won't get ore simplified without going into decimals than x=log(17)/log(3) right?

OpenStudy (anonymous):

*more not ore sorry

jimthompson5910 (jim_thompson5910):

yeah basically \[\Large x = \log_{3}(17)\] and you can use the change of base formula to say \[\Large x = \log_{3}(17) = \frac{\log(17)}{\log(3)}\]

OpenStudy (anonymous):

Ok. thanks again. I really need to get myself together for this HW.

jimthompson5910 (jim_thompson5910):

you're doing quite well actually

OpenStudy (anonymous):

thanks... :P you're a really great helper too. :) I think I can cover the rest of these now that I got what you meant. thank you so much.

jimthompson5910 (jim_thompson5910):

you're welcome